OpenStudy (anonymous):

Assume that 4 digits are selected at random from the set { 1, 2, 4, 6, 7, 9 } and are arranged in random order. What is the probability that the resulting 4-digit number is less than 7000?

OpenStudy (anonymous):

lets say Event A= picking up 4 nos at rando from 6 which is done in 6C4 ways

OpenStudy (anonymous):

event B= number is greater than 7000 ( yes we do this and then we can find its probability, and subtract the same from 1).

OpenStudy (anonymous):

so if we want a number larger than 7000 we should have first digit as 7, or 9 lets look at 4 digit numbers beginnign with 7 fix 7 in first place, then second place you can have other 5 numbers, and in thired place other 4 and in fourth place other 3 nos. similarly do it for 9 and then add these two events and find prob

OpenStudy (anonymous):

So... 4*5*4*3=240 4*5*4*3=240 4*5*4*3=240 4*5*4*3=240 ______________ 240*4?

OpenStudy (anonymous):

nope 5*4*3=60 , for numbers with 7 and 60 for numbers with 9, makes 120

OpenStudy (anonymous):

are all numbers beggining >7000

OpenStudy (anonymous):

I'm still kinda confused

OpenStudy (anonymous):

it is simple. you must use your permutations formula.

OpenStudy (anonymous):

I know how to use them, but I'm not sure how to use them here

OpenStudy (anonymous):

okgive me a minte to work it out.

OpenStudy (anonymous):

you have 720 possiple answers. but you cant use 7 or 9 to start the number. so if you take out all the options with 7 or 9 at the start, then you would do 720-x=y where x is the number of permutation starting with 9 and 7.

OpenStudy (anonymous):

So the ones that start with 9 and 7 would be: 2*5*4*3?

OpenStudy (anonymous):

wait i have that wrong it is a total of 360 answers

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

permutation formula would be 6^P^4 wich makes you have 6x5x4x3.

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

Ahh! How would I find the ones that start with 7 and 9?!

OpenStudy (anonymous):

when is this due?

OpenStudy (anonymous):

Got it! 6P4 - 2*5P3=240 240/360=answer